if 1/3 and 1/4 are the lengths of the two legs of a right triangle, find the length of the hypotenuse
Answer by stanbon(75887)

if 1/3 and 1/4 are the lengths of the two legs of a right triangle, find the length of the hypotenuse
-------
h^2 = (1/3)^2 + (1/4)^2
---
h^2 = 1/9 + 1/16
h^2 = (25)/(9*16)
h = 5/(3*4)
h = 5/12
===============
Cheers,
stan H.